# Last Iterate Convergence of Incremental Methods and Applications in Continual Learning

### v1

@@ 3.2 Convex Lipschitz Setting
=\;= $\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1,t}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k-1,t+1}-{\bm{z}}\|^{2}\big{)}-\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}.$ (3.4)
Combining (3.3) and (3.4) and noticing that ${\bm{x}}_{k-1,T+1}={\bm{x}}_{k}$ and ${\bm{x}}_{k-1,1}={\bm{x}}_{k-1}$, we sum the inequalities over $t\in[T]$ and obtain
@@
We then use the sequence $\{{\bm{z}}_{k}\}$ defined in (2.5) to deduce the last iterate rate in the following theorem. In contrast to (2.1) and (3.1), we do not have the additional error term $f({\bm{z}})-f({\bm{x}}_{*})$ on the right-hand side of (3.2), so the analysis is much simplified.
Theorem 3.4 (IP).
Under Assumptions 1 and 2, the output ${\bm{x}}_{K}$ of Algorithm 2 satisfies
Moreover, given $\epsilon>0$, there exists a constant step size $\eta$ such that $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$ after $\widetilde{\mathcal{O}}\big{(}\frac{G^{2}T\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{2}}\big{)}$ individual gradient evaluations.
@@
Plugging ${\bm{z}}_{k-1}$ defined in (2.5) into (3.2) and multiplying $\eta_{k}w_{k-1}$ on both sides, we obtain
@@
where we also recall that ${\bm{z}}_{-1}={\bm{x}}_{*}$. Unrolling the terms on the left-hand side as (2.9) and choosing $\lambda_{0}=1$, we obtain
$T\eta_{K}w_{K-1}\big{(}f({\bm{x}}_{K})-f({\bm{x}}_{*})\big{)}$ (3.5)
@@
For simplicity and without loss of generality, we make both inequalities tight and choose $w_{k}=\frac{\sum_{j=k}^{K}\eta_{j}}{\sum_{j=k+1}^{K}\eta_{j}}w_{k-1}$. In particular, we choose $w_{k}=\frac{\eta_{K}}{\sum_{j=k+1}^{K}\eta_{j}}$ for $10\leq k\leq K-$ 1 such that $w_{K-1}=1$, then we divide $T\eta_{K}$ on both sides of (3.5) and obtain
@@
We note that the last iterate rate we obtained in Theorem 3.4 matches the prior best known results for the average iterate guarantees for incremental proximal methods (Bertsekas, 2011; Li et al., 2019) in nonsmooth settings, up to a logarithmic factor. Further, we take the $\Theta(\frac{1}{\sqrt{K}})$ step size in the proof only for analytical simplicity, while the diminishing step sizes $\eta_{k}=\Theta(\frac{1}{\sqrt{k}})$ will yield the same rate via a similar analysis, which we omit for brevity.
3.3 Inexact Proximal Point Evaluations
In the last two subsections, we derived our results assuming that the proximal point operator can be evaluated exactly. However, computing the proximal point corresponds to solving a strongly convex problem, which is generally possible to do only up to finite accuracy. When the component functions are smooth, the subproblems resulting from the proximal point evaluations are both smooth and strongly convex and as such can be solved efficiently to high accuracy (in ${\mathcal{O}}(\log(1/\epsilon))$ iterations to reach an error $\epsilon$). For brevity, we thus focus on the nonsmooth convex Lipschitz setting and discuss how to extend our results to admit certain inexactness arisen from approximating the proximal point. In particular, we use the inexactness assumption from Lemaire (1992); Cominetti (1997); Alber et al. (1997); Salzo and Villa (2012); Villa et al. (2013), i.e., where ${\bm{x}}_{k-1,t+1}$ is an approximation of $\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})$ with error $\varepsilon_{k-1,t}>0$ if
$\frac{1}{\eta_{k}}({\bm{x}}_{k-1,t}-{\bm{x}}_{k-1,t+1})\in\partial_{\frac{\varepsilon_{k-1,t}^{2}}{2\eta_{k}}}f_{t}({\bm{x}}_{k-1,t+1}).$ (3.6)
By the results in Salzo and Villa (2012, Proposition 1), the inexact proximal point ${\bm{x}}_{k-1,t+1}$ also satisfies $f_{t}({\bm{x}}_{k-1,t+1})+\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}\leq\min_{{\bm{x}}\in\mathbb{R}^{d}}\Big{\{}f_{t}({\bm{x}})+\frac{1}{2\eta_{k}}\|{\bm{x}}-{\bm{x}}_{k-1,t}\|^{2}\Big{\}}+\frac{\varepsilon_{k-1,t}^{2}}{2\eta_{k}}$. Further using the strong convexity of $f_{t}({\bm{x}})+\frac{1}{2\eta_{k}}\|{\bm{x}}-{\bm{x}}_{k-1,t}\|^{2}$, we have
$\|{\bm{x}}_{k-1,t+1}-\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})\|\leq\varepsilon_{k-1,t}.$
Denoting ${\bm{g}}_{k-1,t}:=\frac{1}{\eta_{k}}({\bm{x}}_{k-1,t}-{\bm{x}}_{k-1,t+1})$ and noticing that $\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})=\frac{1}{\eta_{k}}({\bm{x}}_{k-1,t}-\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t}))$, we have
$\|{\bm{g}}_{k-1,t}-\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})\|\leq\frac{\varepsilon_{k-1,t}}{\eta_{k}}.$ (3.7)
With these insights, we are ready to bound $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ as follows.
Lemma 3.5.
Under Assumptions 1 and 2, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Algorithm 2, we have for $k\in[K]$
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}+\frac{T(T-1)G^{2}\eta_{k}}{2}$ (3.8)
$+\frac{1}{2\eta_{k}}\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}+GT\sum_{t=1}^{T}\varepsilon_{k-1,t}.$
@@ 3.3 Inexact Proximal Point Evaluations
Since $f_{t}$ is convex and closed, we have
$\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})=\frac{1}{\eta_{k}}({\bm{x}}_{k-1,t}-\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t}))\in\partial f_{t}(\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})).$
@@
Decomposing ${\bm{g}}_{k-1,s}={\bm{g}}_{k-1,s}-\nabla M_{\eta_{k}f_{s}}({\bm{x}}_{k-1,s})+\nabla M_{\eta_{k}f_{s}}({\bm{x}}_{k-1,s})$ and using triangle inequality, we have
@@
$\overset{(i)}{\leq}\;$ $\eta_{k}G\sum_{s=t+1}^{T}\Big{(}\frac{\varepsilon_{k-1,s}}{\eta_{k}}+G\Big{)}\leq(T-t)G^{2}\eta_{k}+G\sum_{s=t+1}^{T}\varepsilon_{k-1,s},$ (3.9)
where we use (3.7) and 2 for $(i)$. On the other hand, by the definition of $\frac{\epsilon_{k-1,t}^{2}}{2\eta_{k}}$-subdifferential of $f_{t}$ and using (3.6), we obtain
$f_{t}({\bm{z}})\geq f_{t}({\bm{x}}_{k-1,t+1})+\left\langle{\bm{g}}_{k-1,t},{\bm{z}}-{\bm{x}}_{k-1,t+1}\right\rangle-\frac{\varepsilon_{k-1,t}^{2}}{2\eta_{k}}.$
Expanding the inner product in the above quantity leads to

### v2

@@ B.2 Convex Lipschitz Setting
=\;= $\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1,t}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k-1,t+1}-{\bm{z}}\|^{2}\big{)}-\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}.$ (B.6)
Combining Eq. (B.5) and (B.6) and noticing that ${\bm{x}}_{k-1,T+1}={\bm{x}}_{k}$ and ${\bm{x}}_{k-1,1}={\bm{x}}_{k-1}$, we sum the inequalities over $t\in[T]$ and obtain
@@
See 3.3
@@
Plugging ${\bm{z}}_{k-1}$ defined in Eq. (2.1) into Eq. (B.4) and multiplying $\eta_{k}w_{k-1}$ on both sides, we obtain
@@
where we also recall that ${\bm{z}}_{-1}={\bm{x}}_{*}$. Unrolling the terms on the left-hand side as Eq. (A.5) and choosing $\lambda_{0}=1$, we obtain
$T\eta_{K}w_{K-1}\big{(}f({\bm{x}}_{K})-f({\bm{x}}_{*})\big{)}$ (B.7)
@@
For simplicity and without loss of generality, we make both inequalities tight and choose $w_{k}=\frac{\sum_{j=k}^{K}\eta_{j}}{\sum_{j=k+1}^{K}\eta_{j}}w_{k-1}$. In particular, we choose $w_{k}=\frac{\eta_{K}}{\sum_{j=k+1}^{K}\eta_{j}}$ for $10\leq k\leq K-$ 1 such that $w_{K-1}=1$, then we divide $T\eta_{K}$ on both sides of Eq. (B.7) and obtain
@@
B.3 Inexact Proximal Point Evaluations
We first prove the convergence results for convex smooth settings. The following techical lemma bounds $f({\bm{x}}_{k})-f({\bm{z}})$ within each epoch with inexact proximal point evaluations.
Lemma B.3.
Under Assumptions 1 and 3, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Alg. 2 and for $\alpha>0,\beta>0$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the step sizes satisfy $\eta_{k}\leq\frac{1}{\sqrt{\beta}TL}$, then we have for $k\in[K]$
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $2\eta_{k}^{2}L\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}+\frac{\alpha}{\beta}T\big{(}f({\bm{z}})-f({\bm{x}}_{*})\big{)}+\frac{T}{\eta_{k}}\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}$
$+\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\frac{1}{2\eta_{k}}\big{(}1-\frac{\sum_{t=1}^{T}\varepsilon_{k-1,t}}{\sqrt{\eta_{k}}}\big{)}\|{\bm{x}}_{k}-{\bm{z}}\|^{2}+\frac{\sum_{t=1}^{T}\varepsilon_{k-1,t}}{2\sqrt{\eta_{k}}}.$
@@ B.3 Inexact Proximal Point Evaluations
Since each $f_{t}$ is convex and $L$-smooth, we have for $t\in[T]$
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{x}}_{k-1,t+1})\leq\left\langle\nabla f_{t}({\bm{x}}_{k}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\right\rangle-\frac{1}{2L}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1})\|^{2},$
$f_{t}({\bm{x}}_{k-1,t+1})-f_{t}({\bm{z}})\leq\left\langle\nabla f_{t}({\bm{x}}_{k-1,t+1}),{\bm{x}}_{k-1,t+1}-{\bm{z}}\right\rangle-\frac{1}{2L}\|\nabla f_{t}({\bm{x}}_{k-1,t+1})-\nabla f_{t}({\bm{z}})\|^{2}.$
Following the proof of Lemma 2.1, we add and subtract $\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t}-{\bm{z}}\|^{2}$ on the right-hand side of the second inequality and notice that $\left\langle{\bm{g}}_{k-1,t},{\bm{z}}\right\rangle+\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t}-{\bm{z}}\|^{2}\geq\left\langle{\bm{g}}_{k-1,t},{\bm{x}}_{k-1,t+1}\right\rangle+\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t}-{\bm{x}}_{k-1,t+1}\|^{2}+\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t+1}-{\bm{z}}\|^{2},$ then we combine the above inequalities to obtain
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{z}})\leq\;$ $\left\langle\nabla f_{t}({\bm{x}}_{k}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\right\rangle+\left\langle\nabla f_{t}({\bm{x}}_{k-1,t+1})-{\bm{g}}_{k-1,t},{\bm{x}}_{k-1,t+1}-{\bm{z}}\right\rangle$
We decompose $\nabla f_{t}({\bm{x}}_{k})=\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1})+\nabla f_{t}({\bm{x}}_{k-1,t+1})-{\bm{g}}_{k-1,t}+{\bm{g}}_{k-1,t}$ in the first inner product term on the right-hand side, and sum the inequalities over $t\in[T]$ with noticing ${\bm{x}}_{k-1}={\bm{x}}_{k-1,1}$ and ${\bm{x}}_{k}={\bm{x}}_{k-1,T+1}$, and obtain
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\underbrace{\sum_{t=1}^{T}\left\langle{\bm{g}}_{k-1,t},{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\right\rangle-\frac{1}{2\eta_{k}}\sum_{t=1}^{T}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}}_{{\mathcal{T}}_{1}}$
$+\underbrace{\sum_{t=1}^{T}\left\langle\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\right\rangle}_{{\mathcal{T}}_{2}}$
$+\underbrace{\sum_{t=1}^{T}\left\langle\nabla f_{t}({\bm{x}}_{k-1,t+1})-{\bm{g}}_{k-1,t},{\bm{x}}_{k}-{\bm{z}}\right\rangle}_{{\mathcal{T}}_{3}}+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}$
$-\frac{1}{2L}\sum_{t=1}^{T}\Big{(}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1})\|^{2}+\|\nabla f_{t}({\bm{x}}_{k-1,t+1})-\nabla f_{t}({\bm{z}})\|^{2}\Big{)}.$
${\mathcal{T}}_{1}=-\frac{1}{2\eta_{k}}\|{\bm{x}}_{k}-{\bm{x}}_{k-1}\|^{2}\leq 0.$
For the term ${\mathcal{T}}_{2}$, noticing that ${\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}=-\eta_{k}\sum_{s=t+1}^{T}{\bm{g}}_{k-1,s}$ for $11\leq t\leq T-$ 1 and ${\bm{g}}_{k-1,s}={\bm{g}}_{k-1,s}-\nabla f_{s}({\bm{x}}_{k-1,s+1})+\nabla f_{s}({\bm{x}}_{k-1,s+1})-\nabla f_{s}({\bm{z}})+\nabla f_{s}({\bm{z}})-\nabla f_{s}({\bm{x}}_{*})+\nabla f_{s}({\bm{x}}_{*})$, we use Young’s inequality with parameters $\alpha>0$ and $\beta>0$ to obtain
${\mathcal{T}}_{2}=\;$ $\sum_{t=1}^{T-1}\Big{\langle}\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1}),-\eta_{k}\sum_{s=t+1}^{T}{\bm{g}}_{k-1,s}\Big{\rangle}$
[…]
$+\frac{\alpha\eta_{k}^{2}L}{2}\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\big{(}\nabla f_{s}({\bm{z}})-\nabla f_{s}({\bm{x}}_{*})\big{)}\Big{\|}^{2}+2\eta_{k}^{2}L\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}$
$+\frac{\beta\eta_{k}^{2}L}{2}\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\big{(}\nabla f_{s}({\bm{x}}_{k-1,s+1})-\nabla f_{s}({\bm{z}})\big{)}\Big{\|}^{2}$
$+2\eta_{k}^{2}L\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\big{(}{\bm{g}}_{k-1,s}-\nabla f_{s}({\bm{x}}_{k-1,s+1})\big{)}\Big{\|}^{2}$
For the term ${\mathcal{T}}_{3}$, we use Cauchy-Schwarz inequality and Young’s inequality to get
$\sum_{t=1}^{T}\left\langle\nabla f_{t}({\bm{x}}_{k-1,t+1})-{\bm{g}}_{k-1,t},{\bm{x}}_{k}-{\bm{z}}\right\rangle$
[…]
Further using the fact that $\|\sum_{i=1}^{n}{\bm{x}}_{i}\|^{2}\leq n\sum_{i=1}^{n}\|{\bm{x}}_{i}\|^{2}$ and combining the above bounds on ${\mathcal{T}}_{1}$, ${\mathcal{T}}_{2}$ and ${\mathcal{T}}_{3}$ with $\|{\bm{g}}_{k-1,t}-\nabla f_{t}({\bm{x}}_{k-1,t+1})\|\leq\frac{\varepsilon_{k-1,t}}{\eta_{k}}$ for $t\in[T]$, we obtain
$+\frac{\alpha\eta_{k}^{2}T^{2}L}{2}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{z}})-\nabla f_{t}({\bm{x}}_{*})\|^{2}+2\eta_{k}^{2}L\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}$
$+2T^{2}L\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}+\frac{1}{2\eta_{k}^{3/2}}\sum_{t=1}^{T}\varepsilon_{k-1,t}\|{\bm{x}}_{k}-{\bm{z}}\|^{2}+\frac{1}{2\sqrt{\eta_{k}}}\sum_{t=1}^{T}\varepsilon_{k-1,t}$
It remains to follow the proof of Lemma 2.1 and use $\eta_{k}\leq\frac{1}{\sqrt{\beta}TL}\leq\frac{1}{2TL}$ for $\beta\geq 4$ to obtain
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $2\eta_{k}^{2}L\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}+\frac{\alpha}{\beta}T\big{(}f({\bm{z}})-f({\bm{x}}_{*})\big{)}+\frac{T}{\eta_{k}}\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}$
$+\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\frac{1}{2\eta_{k}}\big{(}1-\frac{\sum_{t=1}^{T}\varepsilon_{k-1,t}}{\sqrt{\eta_{k}}}\big{)}\|{\bm{x}}_{k}-{\bm{z}}\|^{2}+\frac{\sum_{t=1}^{T}\varepsilon_{k-1,t}}{2\sqrt{\eta_{k}}},$
thus finishing the proof. ∎
See 3.4
Proof.
Using Lemma B.3 and following the proof of Theorem 2.3 with multiplying $\eta_{k}w_{k-1}$ on both sides, we have
$T\eta_{k}w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{z}}_{k-1})\big{)}$
[…]
$+\frac{\lambda_{k-1}^{2}w_{k-1}}{2}\|{\bm{x}}_{k-1}-{\bm{z}}_{k-2}\|^{2}-\frac{w_{k-1}(1-\sum_{t=1}^{T}\varepsilon_{k-1,t}/\sqrt{\eta_{k}})}{2}\|{\bm{x}}_{k}-{\bm{z}}_{k-1}\|^{2}.$
Then we sum the above inequality over $k\in[K]$ and follow the proof of Theorem 2.3. To telescope the terms $\|{\bm{x}}_{k}-{\bm{z}}_{k-1}\|^{2}$, we need $\sum_{t=1}^{T}\varepsilon_{k-1,t}/\sqrt{\eta_{k}}\leq 1-\lambda_{k}$ for $11\leq k\leq K-$ 1 such that
$\lambda_{k}^{2}w_{k}\leq\lambda_{k}w_{k}\Big{(}1-\sum_{t=1}^{T}\varepsilon_{k-1,t}/\sqrt{\eta_{k}}\Big{)}\leq w_{k-1}\Big{(}1-\sum_{t=1}^{T}\varepsilon_{k-1,t}/\sqrt{\eta_{k}}\Big{)}.$
In this case, we maintain the same requirements on $\{\lambda_{k}\}$ and $\{w_{k}\}$ to obtain the guarantee on the last iterate as in Theorem 2.3. In particular, we take the same choices with constant step sizes $\eta_{k}\equiv\eta$ such that $\lambda_{k}=\frac{w_{k-1}}{w_{k}}=\frac{(1+\frac{\alpha}{\beta})(K-k)}{1+(1+\frac{\alpha}{\beta})(K-k)}$ for $10\leq k\leq K-$ 1, so it suffices to let $\sum_{t=1}^{T}\varepsilon_{k-1,t}\leq\frac{\sqrt{\eta}}{1+(1+\frac{\alpha}{\beta})(K-k+1)}$ for $1\leq k\leq K$. Following the proof of Theorem 2.3, we obtain
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{w_{-1}}{2\eta T}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}+2\eta^{2}T^{2}\sigma_{*}^{2}L\sum_{k=1}^{K}w_{k-1}$
$+\frac{1}{\eta}\sum_{k=1}^{K}\sum_{t=1}^{T}w_{k-1}\varepsilon_{k-1,t}^{2}+\frac{1}{2\sqrt{\eta}T}\sum_{k=1}^{K}\sum_{t=1}^{T}w_{k-1}\varepsilon_{k-1,t}.$
Plugging in the choice that $w_{k-1}\leq\frac{\mathrm{e}}{(K-k+1)^{\frac{1}{1+\alpha/\beta}}}$ for $11\leq k\leq K-$ 1 and $w_{K-1}=1$, we then have
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{\mathrm{e}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2\eta TK^{\frac{1}{1+\alpha/\beta}}}+2\eta^{2}T^{2}\sigma_{*}^{2}L(1+\beta/\alpha)K^{\frac{\alpha/\beta}{1+\alpha/\beta}}+\frac{\mathrm{e}}{2\eta T}\sum_{k=0}^{K-1}\sum_{t=1}^{T}\frac{2T\varepsilon_{k,t}^{2}+\sqrt{\eta}\varepsilon_{k,t}}{(K-k)^{\frac{1}{1+\alpha/\beta}}}.$
Hence, given $\varepsilon>0$, to maintain the convergence rate with exact proximal point evaluations, it suffices to take $\sum_{t=1}^{T}\varepsilon_{k-1,t}\leq\sqrt{\eta}\min\{\frac{\varepsilon}{4\mathrm{e}^{2}(1+\log K)},\frac{1}{1+(1+\frac{\alpha}{\beta})(K-k+1)}\}$ for $1\leq k\leq K$. Indeed, we have
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{\mathrm{e}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2\eta TK^{\frac{1}{1+\alpha/\beta}}}+2\eta^{2}T^{2}\sigma_{*}^{2}L(1+\beta/\alpha)K^{\frac{\alpha/\beta}{1+\alpha/\beta}}+\sum_{k=0}^{K}\frac{2\mathrm{e}\varepsilon}{(K-k)^{\frac{1}{1+\alpha/\beta}}}$
$\overset{(i)}{\leq}\;$ $\frac{\mathrm{e}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2\eta TK^{\frac{1}{1+\alpha/\beta}}}+2\eta^{2}T^{2}\sigma_{*}^{2}L(1+\beta/\alpha)K^{\frac{\alpha/\beta}{1+\alpha/\beta}}+2\mathrm{e}\varepsilon(1+\beta/\alpha)K^{\frac{\alpha/\beta}{1+\alpha/\beta}}.$
It remains to follow the proof of Theorem 2.3, and we choose $\sum_{t=1}^{T}\varepsilon_{k-1,t}=\sqrt{\eta}\min\{\varepsilon,\frac{1}{3(K-k+1)}\}$, assuming without loss of generality that $\varepsilon\leq\frac{1}{4\mathrm{e}^{2}(1+\log K)}$. ∎
We then come to prove the convergence with inexact proximal point evaluations for convex Lipschitz settings.
Lemma B.4.
Under Assumptions 1 and 2, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Alg. 2, we have for $k\in[K]$
$T(f({\bm{x}}_{k})-f({\bm{z}}))\leq$ $\frac{1}{2\eta_{k}}\Big{(}1+\frac{1}{2\eta_{k}GT}\sum_{t=1}^{T}\varepsilon_{k-1,t}\Big{)}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\frac{1}{2\eta_{k}}\|{\bm{x}}_{k}-{\bm{z}}\|^{2}$ (B.8)
$+\frac{T(T-1)G^{2}\eta_{k}}{2}+\frac{1}{\eta_{k}}\Big{(}\sum_{t=1}^{T}\varepsilon_{k-1,t}\Big{)}^{2}+3GT\sum_{t=1}^{T}\varepsilon_{k-1,t}.$
Proof.
@@
Decomposing ${\bm{g}}_{k-1,s}={\bm{g}}_{k-1,s}-\nabla M_{\eta_{k}f_{s}}({\bm{x}}_{k-1,s})+\nabla M_{\eta_{k}f_{s}}({\bm{x}}_{k-1,s})$ and using triangle inequalities, we have
@@
$\overset{(i)}{\leq}\;$ $\eta_{k}G\sum_{s=t+1}^{T}\Big{(}\frac{\varepsilon_{k-1,s}}{\eta_{k}}+G\Big{)}\leq(T-t)G^{2}\eta_{k}+G\sum_{s=t+1}^{T}\varepsilon_{k-1,s},$ (B.9)
where we use Eq. (3.2) and the fact that $\nabla M_{\eta_{k}f_{s}}({\bm{x}}_{k-1,s})\in\partial f_{t}(\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t}))$ for $(i)$. On the other hand, using convexity of $f_{t}$, we have for $t\in[T]$ that
$f_{t}({\bm{z}})\geq\;$ $f_{t}({\bm{x}}_{k-1,t+1})+\left\langle\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t}),{\bm{z}}-{\bm{x}}_{k-1,t+1}\right\rangle$
=\;= $f_{t}({\bm{x}}_{k-1,t+1})+\left\langle{\bm{g}}_{k-1,t},{\bm{z}}-{\bm{x}}_{k-1,t+1}\right\rangle+\left\langle\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})-{\bm{g}}_{k-1,t},{\bm{z}}-{\bm{x}}_{k-1,t+1}\right\rangle.$
Expanding the inner product in the above quantity and using Cauchy-Schwarz inequality with Eq. (3.2) leads to